Smith engineering is in the process of reviewing the salaris of their surveyors. During this review th company foudn that and experience surveyor can survery a roadbed in 3 hours. An apprentice surveyor needs 7 hours to survey the same stretch of road.
If the two worked together find how loing it takes them to coplete the job

Answer by stanbon(57342)

Smith engineering is in the process of reviewing the salaris of their surveyors. During this review the company foud that and experience surveyor can survey a roadbed in 3 hours. An apprentice surveyor needs 7 hours to survey the same stretch of road.
If the two worked together find how long it takes them to complete the job.
-------------------------
Experience surveyor rate : 1/3 job/hr
Apprentice rate: 1/7 job/hr
Together rate: 1/x job/hr
--------------------------------
Equation:
rate + rate + together rate:
1/3 + 1/7 = 1/x
7x + 3x = 21
10x = 21
x = 2.1 hours (time it will take if they work together)
==========================================================
Cheers,
Stan H.